Part C In terms of mass, calculate the percentage of the active ingredient for each of the following medications. Afterward, rank the percentages from the smallest to highest. Medication Pill mass Active ingredient mass 1.65x10 2 g 25.9 mg Pill A 9.60x10 2 g 478 mg Pill B 1.94x105 Aug 263 mg Pill C 5.75x104 Aug 165 mg Pill D Rank from lowest to highest percent of active ingredient. To rank items as equivalent, overlap them. Reset Help Lowest percent active ingredient Highest percent active ingredient Pill C Pill B Pill D Pill A
Answer #1

Percentage of active ingredient = (mass of active ingredient / Pill mass) x 100

Make sure all the values are in same unit = grams

So after converting them ,

Pill A = (1.65 x 10-2 / 25.9 x 10-3) x 100 = 63.7%

Pill B = (9.60 x 10-2 / 478 x 10-3) x 100 = 20.08%

Pill C = (1.94 x 10-1/ 263 x 10-3) x 100 = 73.8%

Pill D = (5.75 x 10-2/ 165 x 10-3) x 100= 34.8%

so the order from lowest to highest percentage is

Pill B < Pill D < Pill A < Pill C